OpenStudy (anonymous):

int_{0}^{1}int_{8y}^{8}e ^{x ^{2}}dxdy

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{8y}^{8}e ^{x ^{2}}dxdy\]

OpenStudy (anonymous):

 The region is the triangle in the first quadrant indicated above. It's tough to integrate \[e ^{x ^{2}}\] with respect to x (in fact, you can't really do it). So, your best bet is to change the order of integration from dxdy to dydx. If you look at the region I drew, the limits on x are 8y and 8 (as given), so x is bound by x=8y from the left and x=8 from above. The point of intersection of the two lines is (8,1), so the new limits on x would be from 0 to 8. Since x=8y, we know that y=(1/8)x, so the limits on y are 0 to (1/8)x. So our new integral is \[\int\limits_{0}^{8}\int\limits_{0}^{(1/8)x}e ^{x ^{2}}dydx\] Integrating the first part yields \[\int\limits_{0}^{8}(1/8)xe ^{x ^{2}}dx\] If we let \[u=x ^{2}\] Then du=2xdx, xdx=(1/2)du, and (1/8)xdx=(1/16)du Also note that for x=0, u=0 and for x=8, u=64. So our integral becomes \[\int\limits\limits_{0}^{64}(1/16)e ^{u}du\] So we get \[(1/16)e ^{u}\] and evaluate from 0 to 64, yielding \[(1/16)(e ^{64}-1)\] Hope this makes sense.